What is the molecular mass of a non-electrolyte solute in a solution that contains 0.6kg of solute and 200g of benzene with a freezing point equal to 5.4°C? The freezing point of pure benzene is 5.53°C/m. The freezing point constant of benzene is 5.12°C.

1
Expert's answer
2022-04-19T03:11:53-0400

"\Delta T=5.53-5.4=0.13°C=C_mK_f=\frac{m}{M}/m_sK_f=118.12"

"M=\frac{m}{m_s} \frac{K_f}{\Delta T}=\frac{600}{200} \frac{5.12}{0.13}=118.12 g/mol"